[figure: 9.7 cm]

FERROMAGNETIC COUPLING IN QUASI-1D Mn-PORPHYRIN

BASED MOLECULAR MAGNETS

K. Falk^{a}, M. Bałanda^{b}, T. Wasiutyński^{b}, Z. Tomkowicz^{c} and W. Haase^{a} a Darmstadt University of Technology, Petersenstrasse 20, 64287 Darmstadt, Germany

b Institute of Nuclear Physics, 31-342 Kraków, Radzikowskiego 152, Poland cInstitute of Physics, Jagellonian University, Reymonta 4, 30-059 Kraków, Poland

This work concerns magnetism of [MnR_{4}TPP][TCNE] charge-transfer molecular magnets. The initial [MnTPP][TCNE] compound is built of chains with MnTPP cations bound to TCNE anions. Magnetic moments of Mn^{III} and of the anion are strongly coupled in the chain and at T_{c} = 18 K transition from 1D to the bulk magnetic state takes place. The type of magnetic interaction between the chains is not clarified yet. We investigate structural modifications with alongated R = OC_{n}H_{2n+1} (n = 10, 12,

14) groups substituted at the periphery of the porphyrin disc. AC susceptibility and DC magnetization measurements show both ferromagnetic and spin-glass behaviour. In spite of the larger distance between the chains, in comparison to that of [MNTP-P][TCNE], T_{c} is increased up to 22 K. The interchain coupling constant estimated from the quasi-1D model [1] for the compound with n = 12 is ≈ 0.15 K and equals ca. 0.1 % of the intrachain exchange. This value supports the hypothesis [2] according to which, for materials composed of strongly exchange-coupled magnetic chains, the classical dipolar interactions between the chains may stabilize the bulk magnetism.

[1] Shu-Qing Bao, Jue-Lian Shen, Guo-Zhen Yang, Phys. Lett. A 224 (1997) 203.

[2] C.M. Wynn et al., Chem.Materials 9 (1997) 2156.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Maria Bałanda

Address for correspondence : Instytut Fizyki Ja̧drowej im. H. Niewodniczańskiego ul. Radzikowskiego 152 31-342 Kraków

Email address : balanda@alf.ifj.edu.pl